# GVE-Louvain: Fast Louvain Algorithm for Community Detection in Shared Memory Setting

### v2

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents an optimized parallel implementation of Louvain, a high quality community detection method, for shared memory multicore systems. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Louvain, which we term as GVE-Louvain, outperforms Vite, Grappolo, and NetworKit Louvain by 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × respectively - achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. In addition, GVE-Louvain improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
Community detection is the problem of uncovering the underlying structure of complex networks, i.e., identifying groups of vertices that exhibit dense internal connections but sparse connections with the rest of the network, in an unsupervised manner. It is an NP-hard problem with numerous applications in domains such as drug discovery, protein annotation, topic discovery, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). The Louvain method (Blondel et al., 2008) is a popular heuristic-based approach for community detection, with the modularity metric (Newman, 2006) being used to measure the quality of communities identified.
In recent years, the collection of data and the relationships among them, represented as graphs, have reached unmatched levels. This has necessitated the design of efficient parallel algorithms for community detection on large networks. The multicore/shared memory setting is crucial for community detection due to its energy efficiency and the prevalence of hardware with extensive DRAM sizes. However, many of the current algorithms for community detection are challenging to parallelize due to their irregular and inherently sequential nature (Halappanavar et al., 2017), in addition to the complexities of handling concurrency, optimizing data access, reducing contention, minimizing load imbalance. Existing studies on Louvain propose algorithmic optimizations and parallelization techniques, but do not address optimization for the aggregation phase of the Louvain algorithm, which emerges as a bottleneck after the local-moving phase of the algorithm has been optimized. Moreover, these optimization techniques are scattered over a number of papers, making it difficult for a reader to get a grip over them.
1.1. Our Contributions
This report introduces GVE-Louvain, an optimized parallel implementation of Louvain [^1] https://github.com/puzzlef/louvain-communities-openmp for shared memory multicores. On a machine with two 16-core Intel Xeon Gold 6226R processors, GVE-Louvain outperforms Vite, Grappolo, and NetworKit Louvain by 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × respectively, achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. With doubling of threads, GVE-Louvain exhibits an average performance scaling of 1.6×1.6$\times$1.6 ×.
@@ 2. Related work
A few open source implementations and software packages have been developed for community detection. Vite (Ghosh et al., 2018a) is a distributed memory parallel implementation of the Louvain method that incorporates several heuristics to enhance performance while maintaining solution quality, while Grappolo (Halappanavar et al., 2017) is a shared memory parallel implementation. NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes a parallel implementation of the Louvain algorithm.
@@ 4.1.6. Vertex pruning
Vertex pruning is a technique that is used to minimize unnecessary computation. Here, when a vertex changes its community, its marks its neighbors to be processed. Once a vertex has been processed, it is marked as not to be processed. However, it comes with an added overhead of marking an unmarking of vertices. Based on our results, vertex pruning justifies this overhead, and should be enabled for $11\%$ improvement in performance. An illustration of vertex pruning optimization is shown in Figure 1.
@@ 4.1.9. Hashtable design for local-moving/aggregation phases
One can use C++’s inbuilt maps as per-thread (independent) hashtables for the Louvain algorithm. But this has poor performance. So we use a key-list and a full-size values array (collision-free) to dramatically improve performance. However, if the memory addresses of the hashtables are nearby (Close-KV), even if each thread uses its own hashtable exclusively, performance is not as high. This is possibly due to false cache-sharing. Alternatively, if we ensure that the memory address of each hashtable are farther away (Far-KV), the performance improves. Our results indicate that Far-KV has the best performance and is $\times4$ faster than Map, and 1.3×1.3$\times$1.3 × faster than Close-KV hashtable implementations. An illustration of Far-KV hashtable is shown in Figure 2.
@@ 4.2.2. Local-moving phase of GVE-Louvain
Lines 11-23 represent the main loop of the local-moving phase. In line 12, we first initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 22-22, we iterate over unprocessed vertices in parallel. For each vertex $i$, we mark $i$ as processed - vertex pruning (line 14), scan communities connected to $i$ excluding self (line 15), determine the best community $*c*$ to move $i$ to (line 17), calculate the delta-modularity of moving $i$ to $*c*$ (line 18), and update the community membership (lines 20-21) of $i$, and mark its neighbors as unprocessed (line 22) if a better community was found. In line 23, we check if the local-moving phase has converged (locally). If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 24, we return the number of iterations performed $l_{i}$.
@@ 5.1. Experimental Setup
Table 1. List of $13$ graphs obtained SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with ***). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities obtained using Louvain algorithm.
@@ 5.2. Comparing Performance of GVE-Louvain
We now compare the performance of GVE-Louvain with Vite (Louvain), Grappolo (Louvain), and NetworKit Louvain. For Vite, we convert the graph datasets to Vite’s binary graph format, run it on a single node with threshold cycling/scaling optimization, and measure the reported average total time. For Grappolo, we measure the run it on the same system, and measure the reported total time. For NetworKit, we use a Python script to invoke PLM (Parallel Louvain Method), and measure the total time reported with getTiming(). For each graph, we measure the runtime of each implementation five times, for averaging. We also record the modularity of communities obtained, as reported by each implementation.
@@ 6. Conclusion
In conclusion, this study addresses the optimization of Louvain method, a high-quality community detection algorithm, in the shared memory setting. We consider 9 different optimizations, which significantly improve the performance of the local-moving and the aggregation phases of the algorithm. On a server with dual 16-core Intel Xeon Gold 6226R processors, comparison with competitive open source implementations (Vite and Grappolo) and packages (NetworKit) show that our optimized implementation of Louvain, which we term as GVE-Louvain, is on average 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × faster than Vite, Grappolo, and NetworKit respectively. In addition, GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite, and $0.6\%$ lower modularity than Grappolo and NetworKit. On a web graph with $3.8$ billion edges, GVE-Louvain identifies communities in $6.8$ seconds, and thus achieves a processing rate of $560$ million edges/s. In addition, GVE-Louvain achieves a strong scaling factor of 1.6×1.6$\times$1.6 × for every doubling of threads. Looking ahead, future work could focus of designing dynamic algorithms for Louvain to accommodate dynamic graphs which evolve over time. This would contribute to interactive updation of community memberships of vertices in real-world scenarios.

### v3

@@ Abstract.
Community detection is the problem of identifying natural divisions in networks. Efficient parallel algorithms for identifying such divisions is critical in a number of applications, where the size of datasets have reached significant scales. This technical report presents one of the most efficient multicore implementations of the Louvain algorithm, a high quality community detection method. On a server equipped with dual 16-core Intel Xeon Gold 6226R processors, our Louvain, which we term as GVE-Louvain, outperforms Vite, Grappolo, and NetworKit Louvain by 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × respectively - achieving a processing rate of $560M$ edges/s on a $3.8B$ edge graph. In addition, GVE-Louvain improves performance at an average rate of 1.6×1.6$\times$1.6 × for every doubling of threads.
@@ 1. Introduction
Community detection is the problem of uncovering the underlying structure of complex networks, i.e., identifying groups of vertices that exhibit dense internal connections but sparse connections with the rest of the network, in an unsupervised manner. It is an NP-hard problem with numerous applications in domains such as drug discovery, protein annotation, topic discovery, anomaly detection, and criminal identification. Communities identified are intrinsic when based on network topology alone, and are disjoint when each vertex belongs to only one community (Gregory, 2010). One of the difficulties in the community detection problem is the lack of apriori knowledge on the number and size distribution of communities (Blondel et al., 2008). The Louvain method (Blondel et al., 2008) is a popular heuristic-based approach for community detection, with the modularity metric (Newman, 2006) being used to measure the quality of communities identified.
In recent years, the collection of data and the relationships among them, represented as graphs, have reached unmatched levels. This has necessitated the design of efficient parallel algorithms for community detection on large networks. Existing studies on Louvain propose several optimizations (Rotta and Noack, 2011; Waltman and Eck, 2013; Gach and Hao, 2014; Traag, 2015; Lu et al., 2015; Ryu and Kim, 2016; Ozaki et al., 2016; Naim et al., 2017; Halappanavar et al., 2017; Ghosh et al., 2018b; Traag et al., 2019; Zhang et al., 2021; Shi et al., 2021; You et al., 2022; Aldabobi et al., 2022) and parallelization techniques (Cheong et al., 2013; Wickramaarachchi et al., 2014; Lu et al., 2015; Zeng and Yu, 2015; Que et al., 2015; Naim et al., 2017; Fazlali et al., 2017; Halappanavar et al., 2017; Zeitz, 2017; Ghosh et al., 2018b; Bhowmik and Vadhiyar, 2019; Gheibi et al., 2020; Shi et al., 2021; Bhowmick et al., 2022). Further, significant research effort has been focused on developing efficient parallel implementations of Louvain algorithm for multicore CPUs (Staudt and Meyerhenke, 2015; Staudt et al., 2016; Fazlali et al., 2017; Halappanavar et al., 2017; Qie et al., 2022), GPUs (Naim et al., 2017), CPU-GPU hybrids (Bhowmik and Vadhiyar, 2019; Mohammadi et al., 2020), multi-GPUs (Cheong et al., 2013; Hricik et al., 2020; Chou and Ghosh, 2022; Gawande et al., 2022), and multi-node systems — CPU only (Ghosh et al., 2018b, a; Sattar and Arifuzzaman, 2022) / CPU-GPU hybrids (Bhowmick et al., 2022).
However, many of the aforementioned works concentrate on optimizing the local-moving phase of the Louvain algorithm, but do not address optimization for the aggregation phase of the algorithm, which emerges as a bottleneck after the local-moving phase has been optimized. These optimization techniques are also scattered over a number of papers, making it difficult for a reader to get a grip over them. Moreover, much attention has been directed towards GPU-based solutions. However, developing algorithms that efficiently utilize GPUs can be challenging both in terms of initial implementation and ongoing maintenance. Further, the soaring prices of GPUs present hurdles. The multicore/shared memory environment holds significance for community detection, owing to its energy efficiency and the prevalence of hardware with ample DRAM capacities. Through our implementation of the Louvain algorithm, we aim to underscore that CPUs remain adept at irregular computation. Additionally, we show that achieving optimal performance necessitates a focus on the data structures.
In this report, we introduce our parallel multicore implementation of the Louvain algorithm [^1] https://github.com/puzzlef/louvain-communities-openmp. Our implementation employs asynchronous computation, utilizes parallel prefix sum and preallocated Compressed Sparse Row (CSR) data structures for identifying community vertices and storing the super-vertex graph during the aggregation phase, uses fast collision-free per-thread hash tables for the local-moving and aggregation phases, and incorporates an aggregation tolerance to avoid unnecessary aggregation phases. Additionally, we leverage established techniques such as OpenMP’s dynamic loop schedule, limiting the number of iterations per pass, threshold-scaling optimization, and vertex pruning. For each optimization, we determine optimal parameter settings, if there are any. To the best of our knowledge, our implementation represents the most efficient parallel Louvain algorithm implementation on multicore CPUs. We compare our implementation against other state-of-the-art implementations, including multi-core, multi-node, hybrid GPU, multi-GPU, and multi-node multi-GPU implementations, in Table 1. Both direct and indirect comparisons are provided, with details given in Sections 5.2 and A.1 respectively.
Table 1. Speedup of our multicore implementation of Louvain algorithm compared to other state-of-the-art implementations. Direct comparisons are based on running the given implementation on our server, while indirect comparisons (denoted with a *, details given in Section A.1) involve comparing results obtained by the given implementation relative to a common reference (Grappolo). Among the Louvain implementations, some are multi-core, while others are categorized as multi-node (nodes), hybrid GPU (GPU), multi-GPU (GPUs, 1 GPU), and multi-node multi-GPU (DGPUs).
Louvain implementation Published Our Speedup
Grappolo (Halappanavar et al., 2017) 2017 $\times2$
Vite (Ghosh et al., 2018a) 2018 50×50$\times$50 ×
NetworKit Louvain (Staudt et al., 2016) 2016 20×20$\times$20 ×
Qie et al. (Qie et al., 2022) 2022 $273\times^{*}$
PLM (Staudt and Meyerhenke, 2015) 2015 $48\times^{*}$
APLM (Fazlali et al., 2017) 2017 $7.7\times^{*}$
DPLAL (nodes) (Sattar and Arifuzzaman, 2022) 2022 $5472\times^{*}$
Ghosh et al. (nodes) (Ghosh et al., 2018b) 2018 $5.9\times^{*}$
HyDetect (GPU) (Bhowmik and Vadhiyar, 2019) 2019 $54\times^{*}$
Rundemanen (GPU) (Naim et al., 2017) 2017 $9.8\times^{*}$
ACLM (GPU) (Mohammadi et al., 2020) 2020 $6.0\times^{*}$
Nido (1 GPU) (Chou and Ghosh, 2022) 2022 $9.2\times^{*}$
cuVite (1 GPU) (Gawande et al., 2022) 2022 $6.7\times^{*}$
cuGraph (1 GPU) (Hricik et al., 2020) 2020 $4.8\times^{*}$
Cheong et al. (GPUs) (Cheong et al., 2013) 2013 $8.3\times^{*}$
Bhowmick et al. (DGPUs) (Bhowmick et al., 2022) 2022 $4.9\times^{*}$
@@ 2. Related work
We now discuss about a number of state-of-the-art implementation of Parallel Louvain. Mohammadi et al. (Mohammadi et al., 2020) propose the Adaptive CUDA Louvain Method (ACLM), employing GPU acceleration. Their approach involves evaluating the change in modularity for each edge in the graph, reflecting the impact of moving the source vertex to the community of the target vertex. Notably, only the computation of modularity changes occurs on the GPU, with other algorithmic steps executed on the CPU. Sattar and Arifuzzaman (Sattar and Arifuzzaman, 2022) discuss their Distributed Parallel Louvain Algorithm with Load-balancing (DPLAL). They begin by partitioning the input graph using METIS, employing the edge-cut partitioning approach. In each iteration, they determine the optimal community to relocate to for each vertex, execute community changes while tackling duality concerns, evaluate the new modularity, and generate the subsequent level graph. This iterative process continues until no further enhancement in modularity is achieved. Nonetheless, they conduct only a single iteration of the local-moving phase per iteration, which might result in the identification of substandard quality communities. Ghosh et al. (Ghosh et al., 2018a) propose Vite, a distributed memory parallel implementation of the Louvain method that incorporates several heuristics to enhance performance while maintaining solution quality, while Grappolo, by Halappanavar et al. (Halappanavar et al., 2017), is a shared memory parallel implementation. Qie et al. (Qie et al., 2022) present a graph partitioning algorithm that divides the graph into sets of partitions, aiming to minimize inter-partition communication delay and avoid community swaps, akin to the graph coloring approach proposed by Halappanavar et al. (Halappanavar et al., 2017). We do not observe the community swap issue on multicore CPUs (it likely resolves itself), but do observe it on GPUs (likely due to lockstep execution).
Bhowmick et al. (Bhowmik and Vadhiyar, 2019) introduce HyDetect, a community detection algorithm designed for hybrid CPU-GPU systems, employing a divide-and-conquer strategy. The algorithm partitions the graph among the CPU and GPU components of a node, facilitating independent community detection utilizing the Louvain algorithm in both segments. Bhowmick et al. (Bhowmick et al., 2022) later introduce a multi-node multi-GPU Louvain community detection algorithm, which involves graph partitioning across multiple nodes, refinement through identification and migration of doubtful vertices between processors, and utilization of a hierarchical merging algorithm ensuring that the merged components can be accommodated within a processor at any given point. Chou and Ghosh (Chou and Ghosh, 2022) present Nido, a batched clustering method for GPUs that leverages a partition-based design, and can process graphs larger than the combined GPU memory of a node. Gawande et al. (Gawande et al., 2022) propose cuVite, an ongoing endeavor focusing on distributed Louvain for heterogeneous systems — building upon their previous work involving parallelizing Louvain for community detection on CPU-only distributed systems. Finally, NetworKit (Staudt et al., 2016) is a software package designed for analyzing the structural aspects of graph data sets with billions of connections. It is implemented as a hybrid with C++ kernels and a Python frontend, and includes a parallel implementation of the Louvain algorithm.
However, most of the existing works focus on optimizing the local-moving phase of the Louvain algorithm but lack effective parallelization. For instance, the implementation of NetworKit Louvain exhibits several shortcomings. It employs plain OpenMP parallelization for certain operations, utilizing a static schedule with a chunk size of 1, which may not be optimal when threads are writing to adjacent memory addresses. Additionally, NetworKit Louvain employs guided scheduling for the local-moving phase, whereas we utilize dynamic scheduling for better performance. NetworKit Louvain also generates a new graph for each coarsening step, leading to repeated memory allocation and preprocessing due to recursive calls. The coarsening involves several sequential operations, and adding each edge to the coarsened graph requires $O(D)$ operations, where $D$ represents the average degree of a vertex, which is suboptimal for parallelism. Lastly, NetworKit Louvain lacks parallelization for flattening the dendrogram, potentially hindering its performance. Our parallel implementation of Louvain addresses these issues.
@@ 4.1.6. Vertex pruning
Vertex pruning is a technique that is used to minimize unnecessary computation (Ryu and Kim, 2016; Ozaki et al., 2016; Zhang et al., 2021; Shi et al., 2021). Here, when a vertex changes its community, its marks its neighbors to be processed. Once a vertex has been processed, it is marked as not to be processed. However, it comes with an added overhead of marking an unmarking of vertices. Based on our results, vertex pruning justifies this overhead, and should be enabled for $11\%$ improvement in performance. An illustration of this is shown in Figure 1.
@@ 4.1.9. Hashtable design for local-moving/aggregation phases
One can use C++’s inbuilt maps as per-thread (independent) hashtables for the Louvain algorithm. But this has poor performance. So we use a key-list and a full-size values array (collision-free) to dramatically improve performance. However, if the memory addresses of the hashtables are nearby (Close-KV) — as with NetworKit Louvain (Staudt et al., 2016), even if each thread uses its own hashtable exclusively, performance is not as high. This is possibly due to false cache-sharing. Alternatively, if we ensure that the memory address of each hashtable are farther away (Far-KV), the performance improves. Our results indicate that Far-KV has the best performance and is $\times4$ faster than Map, and 1.3×1.3$\times$1.3 × faster than Close-KV hashtable implementations. An illustration of Far-KV hashtable is shown in Figure 2.
@@ 4.2.2. Local-moving phase of GVE-Louvain
Lines 11-23 represent the main loop of the local-moving phase. In line 12, we first initialize the total delta-modularity per iteration $\Delta Q$. Next, in lines 22-22, we iterate over unprocessed vertices in parallel. For each vertex $i$, we mark $i$ as processed - vertex pruning (line 14), scan communities connected to $i$ excluding self (line 15), determine the best community $c*$ to move $i$ to (line 17), calculate the delta-modularity of moving $i$ to $c*$ (line 18), and update the community membership (lines 20-21) of $i$, and mark its neighbors as unprocessed (line 22) if a better community was found. In line 23, we check if the local-moving phase has converged (locally). If so, we break out of the loop (or if $MAX\_ITERATIONS$ is reached). At the end, in line 24, we return the number of iterations performed $l_{i}$.
@@ 5.1. Experimental Setup
Table 2. List of $13$ graphs obtained SuiteSparse Matrix Collection (Kolodziej et al., 2019) (directed graphs are marked with *). Here, $|V|$ is the number of vertices, $|E|$ is the number of edges (after adding reverse edges), $D_{avg}$ is the average degree, and $|\Gamma|$ is the number of communities obtained using Louvain algorithm.
@@ 5.2. Comparing Performance of GVE-Louvain
We now compare the performance of GVE-Louvain with Vite (Louvain) (Ghosh et al., 2018a), Grappolo (Louvain) (Halappanavar et al., 2017), and NetworKit Louvain (Staudt et al., 2016). For Vite, we convert the graph datasets to Vite’s binary graph format, run it on a single node with threshold cycling/scaling optimization, and measure the reported average total time. For Grappolo, we measure the run it on the same system, and measure the reported total time. For NetworKit, we use a Python script to invoke PLM (Parallel Louvain Method), and measure the total time reported with getTiming(). For each graph, we measure the runtime of each implementation five times, for averaging. We also record the modularity of communities obtained, as reported by each implementation.
@@ 6. Conclusion
In conclusion, this report presents our parallel multicore implementation of the Louvain algorithm — a high quality community detection method, which, as far as we are aware, stands as the most efficient implementation of the algorithm on multicore CPUs. Here, we considered $9$ different optimizations, which significantly improve the performance of the local-moving and the aggregation phases of the algorithm. On a server with dual 16-core Intel Xeon Gold 6226R processors, comparison with competitive open source implementations (Vite and Grappolo) and packages (NetworKit) show that our optimized implementation of Louvain, which we term as GVE-Louvain, is on average 50×50$\times$50 ×, $\times2$, and 20×20$\times$20 × faster than Vite, Grappolo, and NetworKit respectively. In addition, GVE-Louvain on average obtains $3.1\%$ higher modularity than Vite, and $0.6\%$ lower modularity than Grappolo and NetworKit. On a web graph with $3.8$ billion edges, GVE-Louvain identifies communities in $6.8$ seconds, and thus achieves a processing rate of $560$ million edges/s. In addition, GVE-Louvain achieves a strong scaling factor of 1.6×1.6$\times$1.6 × for every doubling of threads. Looking ahead, future work could focus of designing dynamic algorithms for Louvain to accommodate dynamic graphs which evolve over time.
@@ Acknowledgements.
Appendix A Appendix
A.1. Indirect Comparison with State-of-the-art Louvain Implementations
We now conduct an indirect comparison of the performance of our multicore implementation of the Louvain algorithm with other similar state-of-the-art implementations, as listed in Table 1. Mohammadi et al. (Mohammadi et al., 2020) introduce the Adaptive CUDA Louvain Method (ACLM), which employs GPU acceleration. For a computational setup featuring a $12$-core AMD Opteron 6344 CPU clocked at $2.60$ GHz alongside an NVIDIA Tesla K20Xm GPU, Mohammadi et al. report a runtime of $0.47$ seconds for ACLM when applied to the in-2004 graph with $13.6$ million edges (refer to Table 6 in (Mohammadi et al., 2020)). Moreover, they document runtimes of $3.71$, $0.60$, and $0.95$ seconds for PLM (Staudt and Meyerhenke, 2015), APLM (Fazlali et al., 2017), and Rundemanen (Naim et al., 2017) respectively (also presented in Table 6 of (Mohammadi et al., 2020)). In contrast, leveraging our system equipped with two $16$-core Intel Xeon Gold 6226R CPUs operating at $2.90$ GHz (delivering up to 3.0×3.0$\times$3.0 × higher performance), we process the indochina-2004 graph boasting $341$ million edges (25.1×25.1$\times$25.1 × larger) in merely $0.65$ seconds. Consequently, our Louvain implementation outperforms ACLM, PLM, APLM, and Rundemanen by approximately 6.0×6.0$\times$6.0 ×, 48×48$\times$48 ×, $\times7$, and 12.2×12.2$\times$12.2 × respectively.
Sattar and Arifuzzaman (Sattar and Arifuzzaman, 2022) propose their Distributed Parallel Louvain Algorithm with Load-balancing (DPLAL). On a graph containing $1M$ vertices, utilizing $40$ compute nodes of the Louisiana Optical Network Infrastructure (LONI) QB2 cluster, where each node is equipped with two $10$ core $2.80$ GHz Intel Xeon E5-2680v2 CPUs, they achieve community detection within approximately $80$ seconds (refer to Figure 4 in their paper (Sattar and Arifuzzaman, 2022)). In contrast, employing our system (which is 6.0×6.0$\times$6.0 × slower, assuming DPLAL utilizes each node with only $25\%$ efficiency), we process the indochina-2004 graph with $7.41M$ vertices (7.41×7.41$\times$7.41 × larger) in only $0.65$ seconds. Consequently, our Louvain implementation is roughly 5472×5472$\times$5472 × faster.
Qie et al. (Qie et al., 2022) introduce a graph partitioning algorithm designed to minimize inter-partition communication delay and avoid community swaps. Their experiments on an $8$ core Intel i9 9900K CPU running at $3.60$ GHz, using the com-LiveJournal graph, reveal a runtime of approximately $1050$ seconds (as depicted in Figure 3 of their paper (Qie et al., 2022)). In contrast, on our system (roughly 3.2×3.2$\times$3.2 × faster), we process the same graph in merely $1.2$ seconds. Consequently, our Louvain implementation achieves a speedup of about 273×273$\times$273 ×.
Bhowmick et al. (Bhowmik and Vadhiyar, 2019) present HyDetect, a community detection algorithm tailored for hybrid CPU-GPU systems, employing a divide-and-conquer strategy. Their experimentation on a system comprising a six-core Intel Xeon E5-2620 CPU operating at $2.00$ GHz paired with an NVIDIA Kepler K40M GPU yields a runtime of $1123$ seconds when applied to the it-2004 graph (as documented in Table 5 of their paper (Bhowmik and Vadhiyar, 2019)). In contrast, leveraging our system equipped with two $16$ core Intel Xeon Gold 6226R CPUs running at $2.90$ GHz (approximately $\times7$ faster), we process the same graph in $2.7$ seconds. Consequently, our Louvain implementation achieves a speedup of around 54×54$\times$54 × compared to HyDetect, without using a GPU.
Bhowmick et al. (Bhowmick et al., 2022) later introduce a multi-node multi-GPU Louvain community detection algorithm. Their experimentation conducted on a CrayXC40 system, where each node comprises one $12$ core $2.40$ GHz Intel Xeon Ivybridge E5-2695 v2 CPU and one NVIDIA Tesla K40 GPU, demonstrates the capability to identify communities within approximately $32$ seconds on the uk-2005 graph utilizing $8$ compute nodes (refer to Figure 8 in their paper (Bhowmick et al., 2022)). Furthermore, they observe a speedup of about 1.7×1.7$\times$1.7 × compared to Cheong et al. (Cheong et al., 2013) on the uk-2005 graph (as depicted in Figure 11 of their paper), and a speedup of approximately 1.2×1.2$\times$1.2 × relative to Ghost et al. (Ghosh et al., 2018b) on the sk-2005 graph (as shown in Figure 10 of their paper). In contrast, on our system featuring two $16$ core Intel Xeon Gold 6226R CPUs operating at $2.90$ GHz (up to 1.6×1.6$\times$1.6 × faster, assuming Bhowmick et al. utilize each compute node with only $25\%$ efficiency), we process the uk-2005 graph in a mere $4.1$ seconds. Thus, our Louvain implementation achieves at least 4.9×4.9$\times$4.9 ×, 8.3×8.3$\times$8.3 ×, and 5.9×5.9$\times$5.9 × speedup compared to Multi-node HyDetect (Bhowmick et al., 2022), Cheong et al. (Cheong et al., 2013), and Ghost et al. (Ghosh et al., 2018b), respectively, without the use of a GPU.
Chou and Ghosh (Chou and Ghosh, 2022) introduce Nido, a batched clustering method designed for GPUs. Their experimentation conducted on a system equipped with two $128$ core AMD EPYC 7742 CPUs running at $2.25$ GHz and eight NVIDIA A100 GPUs demonstrates a (geometric) mean speedup of 2.4×2.4$\times$2.4 × compared to Grappolo (Halappanavar et al., 2017), utilizing $128$ threads (as shown in Table 3 of their paper (Chou and Ghosh, 2022)). Additionally, they find that cuGraph (Hricik et al., 2020) offers a mean speedup of $\times7$ relative to Grappolo (also depicted in Table 3 of their paper). In contrast, our Louvain implementation achieves a mean speedup of $\times2$ compared to Grappolo. Consequently, it surpasses Chou and Ghosh by 9.2×9.2$\times$9.2 × and cuGraph by 2.9×2.9$\times$2.9 ×, again without employing a single GPU.
Gawande et al. (Gawande et al., 2022) introduce cuVite, an ongoing project focusing on distributed Louvain for heterogeneous systems. They evaluate cuVite, cuGraph, and Rundemanen on a single-GPU of NVIDIA DGX-2, i.e., NVIDIA Tesla V100, and with two $24$ core Intel Xeon Platinum 8168 CPUs running at $2.7$ GHz. Further, they evaluate Grappolo on a $224$ core system with eight $28$ core Intel Xeon Platinum 8276M CPUs running at $2.20$ GHz. They observe that, on average, cuGraph, Rundemanen, and cuVite exhibit speedups of $\times3$, 3.0×3.0$\times$3.0 ×, and 3.28×3.28$\times$3.28 × respectively compared to Grappolo. Consequently, our Louvain implementation achieves a speedup of 6.7×6.7$\times$6.7 × compared to both cuGraph and cuVite, and a speedup of 7.3×7.3$\times$7.3 × relative to Rundemanen, all achieved without the utilization of a GPU.